A firm has a production function given by f(x1,x2,x3,x4)=min{2x1+x2,x3+2x4}. What is the cost function for this technology?


1
Expert's answer
2021-05-26T13:28:36-0400

Production function is "f(x_1,x_2,x_3,x_4)=min(2x_1+x_2,x_3+2x_4)"

"x_1\space and\space x_2" are perfect substitutes, and "x_3\space and\space x_4" are also perfect substitutes.

If "w_i" denotes the price per unit of input "x_i," then

"x_1"  is used to produce the output when "\frac{w_1}{2}<w_2"   and "x_2=0"

"x_2"  is used to produce the output when "\frac{w_1}{2}<w_2"   and "x_1=0"

"x_3"  is used to produce the output when "w_3<\frac{w_4}{2}"   and "x_4=0"

"x_4"  is used to produce the output when "w_3>\frac{w_4}{2}"   and "x_3=0"

So if one wants to produce one unit of output when the input prices satisfy "\frac{w_1}{2}<w_2" and "w_3<\frac{w_4}{2}" then one must employ "x_1=\frac{1}{2}" and "x_3=1" so as to minimize cost, and the cost is given by "\frac{w_1}{2}+w_3"

When the input prices satisfy "\frac{w_1}{2}>w_2" and "w_3<\frac{w_4}{2}" then one must employ "x_2=1" and "x_2=1" in order to minimize cost, and the cost is given by "w_2+w_3"

For other two combinations.

we can write the cost of producing one unit of output as

"min(\frac{w_1}{2},w_2)+mim(w_3,\frac{w_4}{2})"

More generally, if you want to produce q units, cost function is

"C(w_1,w_2,w_3,w_4,q)=[min(\frac{w_1}{2},w_2)+min(w_3,\frac{w_4}{2})]q"